OpenStudy (anonymous):

[x^((4n^2)+n)]/[x^(n+1)^(n-1)] *problem in comments*

OpenStudy (anonymous):

\[(x^{4n ^{2}+n})\div (x ^{n+1})^{(x-1)}\]

OpenStudy (anonymous):

x^((4n^2)+n)-(n^2-1))=x^(3n^2 +n+1)

OpenStudy (jiteshmeghwal9):

Denominator use the property :- \(\Large{{(x^b)^c=x^{bc}}}\)

OpenStudy (anonymous):

Is it ? \[\large \frac{x^{4n^2 + n}}{(x^{n+1})^{n-1}}\]

OpenStudy (jiteshmeghwal9):

using this property what do u get as denominator @spasta106 ??

OpenStudy (anonymous):

I got x^(n^2-1)

OpenStudy (anonymous):

@waterineyes yes that's the problem

OpenStudy (jiteshmeghwal9):

now use the property :- \(\Large{\frac{x^a}{x^b}=x^{a-b}}\)

OpenStudy (anonymous):

Jitesh is guiding you very well @spasta106 Do what he is saying..

OpenStudy (jiteshmeghwal9):

@spasta106 u will get \[\Huge{x^{(4n^2+n)-(n^2-1)}}\]

OpenStudy (anonymous):

Yes, that's what I have so far

OpenStudy (jiteshmeghwal9):

now solve this powers & u will get your answer :)

OpenStudy (jiteshmeghwal9):

x^? @spasta106

OpenStudy (anonymous):

Ok 3n^2+n+1?

OpenStudy (anonymous):

with x as a base

OpenStudy (anonymous):

I'm still getting the same answer...

OpenStudy (jiteshmeghwal9):

ok ! now that is your answer :)

OpenStudy (anonymous):

Ok! thanks!

OpenStudy (jiteshmeghwal9):

yw :)